# DOCE: Finding the Sweet Spot for Execution-Based Code Generation

### v1

@@ 3.4 Self-Debugging
Figure 3: Performance of MBR-Exec with and without filtering over different numbers of generated candidates using CodeLlama-7B-Instruct. We also provide the oracle with solid lines, denoted as Pass@k. Note that results that end with $+$ mean that it is evaluated on the plus with extended test cases. Results are averaged across at least 4 runs.
4 Results
In this section, we provide results and discuss the effects of reranking and self-debugging. Firstly, we analyze the impact of candidate generation on the model’s oracle performance (assuming we pick the best possible candidate) and the performance of execution-based reranking methods. We then compare different reranking methods and discuss the importance of execution for reranking. With the answers to these questions in mind, we study the effect of self-debugging and discuss the best way to combine all these methods to improve reranking performance. We provide a subset of all our experimental results to answer these research questions. For results on more experiments, please refer to Appendix A.
4.1 Candidate Generation
First, we study the impact of candidate generation on reranking. To do so, we vary the number of candidates sampled from CodeLlama-7B-Instruct, also exploring different temperature parameters. We also provide estimates of the oracle of reranking using Pass@k (Chen et al., 2021; Shi et al., 2022), commonly used as an evaluation metric for code generation models. Pass@k measures the proportion that at least one candidate passes all evaluation unit tests given $k$ generated candidates. We evaluate the performance of reranking on EvalPlus (Liu et al., 2023) for rigorous evaluation.
@@ 4.1 Candidate Generation
Figure 4: Performance of MBR-Exec with and without filtering over sampling temperatures using CodeLlama-7B-Instruct with 50 generated candidates. We also provide the oracle of reranking with dashed lines. Note that results that end with $+$ mean that it is evaluated on the plus with extended test cases. Results are averaged across 4 runs.
What is the impact of the number of candidates?
We generate candidates using CodeLlama-7B-Instruct by setting temperature 1.6 and p 0.95. We hypothesize that filtering based on trial unit tests gives reranking better alignment to the ground-truth function regarding execution results.
Figure 3 shows the performance as a function of the number of generated candidates for different reranking approaches (based on execution) and the oracle performance for both HumanEval and MBPP-S. First of all, when filtering is not included, the performance of MBR-Exec only shows moderate improvement with more generated candidates across different numbers of unit test cases for evaluation and two benchmarks, which aligns with the analysis of Shi et al. (2022). However, with filtering, MBR-Exec shows considerable improvements over the version without filtering, and performance scales much better as more candidates are considered, which aligns with our hypothesis that filtering based on trial unit tests gives reranking better alignment to the ground-truth function regarding execution results.
Moreover, the performance of MBR-Exec with filtering consistently approaches the oracle performance even as we increase the number of candidates, which gives further evidence to our hypothesis that filtering based on trial unit tests gives reranking better alignment to the ground-truth function regarding execution results. Additionally, MBR with filtering has closer-to-oracle performance on the benchmark of MBPP-S tested on different setups with unit tests (see Figure 19(d)), while MBR with filtering approaches the oracle performance on both datasets (see Figure 19(c)), the gap is larger on the HumanEval dataset. This suggests that the overlaps of testing ability between trial unit tests and unit tests used for final evaluation are positively correlated. In contrast, 80% of the trial test cases are either included in the original test cases or the extended test cases from EvalPlus (Liu et al., 2023).
What is the impact of sampling temperature?
We compare generations with temperatures from 0.2 to 2.0 using CodeLlama-7B-Instruct with $p$ 0.95. We show results with 50 generated candidates in Figure 4. Note that previous studies on the optimal temperature are either based on worse-performing language models(Li et al., 2022), a different partition of the dataset (Shi et al., 2022), or a setting with trial unit tests not provided with the input (Chen et al., 2023).
When studying the effect of sampling temperature on the open-sourced CodeLlama-7B-Instruct and including the trial unit test in the prompt, we find that the temperature that gives the higher oracle for reranking is much higher than previously seen, with performance peaking at around 1.6 (while previous studied finding values <1 (Chen et al., 2021; Li et al., 2022; Shi et al., 2022; Chen et al., 2023)). We hypothesize that this results from better models allowing the generation of diverse and high-quality code.
Additionally, we observe a substantial difference in reranking performance with increased temperature when filtering based on the trial unit test included, for which we highlight the importance of filtering when reranking is considered. Finally, we adopt sampling temperature 1.6 with nucleus $p$ 0.95 for later experiments. For DeepSeekCoder-6.7B-Instruct, we use temperature 1.2.666With temperature ¿1.2 on DeepSeekCoder-6.7B-Instruct, sampling is more likely to generate token indices that are not defined in the vocabulary, see htatps://github.com/vllm-project/vllm/pull/3685.
HE (+) MBPP-S (+)
Random 34.1 (30.2) 50.1 (41.9)
Greedy 43.3 (39.0) 53.2 (44.8)
Oracle 82.4 (76.2) 83.2 (72.0)
@@
LL 45.1 (38.7) 54.2 (42.6)
CR 46.0 (40.5) 53.1 (43.1)
CS 33.1 (30.2) -
Filter 70.0 (59.7) 77.4 (60.8)
Filter + LL 71.0 (60.4) 78.1 (59.7)
Filter + CR 71.3 (61.3) 78.5 (60.2)
Filter + CS 71.2 (61.0) -
@@
LL + MBR-Exec 51.8 (49.4) 64.4 (58.1)
CR + MBR-Exec 51.4 (48.9) 64.8 (58.7)
Filter + MBR-CBS 69.8 (60.1) 77.3 (61.3)
Filter + MBR-CS 71.5 (59.9) -
Filter + MBR-Exec 74.5 (70.1) 79.0 (69.3)
Filter + LL + MBR-Exec 74.5 (70.6) 78.9 (69.2)
Filter + CR + MBR-Exec 74.0 (70.0) 79.3 (69.4)
Table 1: Performance across different reranking methods without and with filtering by sampling 50 candidates with temperature 1.6 and $p$ 0.95 using Code-Llama-7B-Instruct. For N-Best Reranking, we compare log-likelihood (LL), Coder-Reviewer (CR) (Zhang et al., 2022), and CodeScore (CS) Dong et al. (2023). For MBR, we compare CodeScore, CodeBertScore (CBS) (Zhou et al., 2023), and MBR-Exec (Shi et al., 2022). Note that HE stands for the HumanEval benchmark, $+$ refers to evaluation on extended test cases on EvalPlus. We also highlighted best and second best reranking results of the class of reranking methods.
@@ 4.2 Comparing Reranking Methods
We conduct experiments with $n$-best reranking and MBR to compare different raking approaches with and without filtering based on trial unit tests. We show our results in Table 1.
Filtering is the most powerful method for $n$-best reranking.
According to the group “N-Best Reranking” (see Table 1), filtering based on trial unit tests gives the largest gain from the random baseline compared to MAP decoding, CoderReviewer, and CodeScore. Note that while CodeScore is trained to predict the execution accuracy of generated code, it does not even outperform the random baseline, suggesting the gap in training execution-free rerankers for execution-based metrics.
Not only is filtering based on trial unit tests powerful in itself, but it also boosts the performance of other $n$-best reranking methods substantially. Specifically, we find that all methods including MAP decoding, CoderReviewer, and CodeScore outperform the filtering baseline moderately when combined with filtering except for the MAP decoding on MBPP-S+. Given our findings, we again emphasize the importance of filtering based on unit tests for reranking, which is a statement that is missing from previous studies that involve reranking (Shi et al., 2022; Zhang et al., 2022; Ni et al., 2023; Chen et al., 2024). Additionally, we suggest training better execution-free metrics to approximate reranking performance with filtering based on trial unit tests.
@@
According to Zhang et al. (2022), Coder-Reviewer outperforms MBR-Exec when filtering means if the candidate can give outputs based on the trial unit test inputs, and when MBR-Exec is only implemented on the trial unit tests. However, in our context based on a more sensible definition of filtering and MBR-Exec utilizing all unit test inputs, we found that both MAP and CoderReviewer underperform MBR-Exec with filtering. Additionally, when combined with filtering and MBR-Exec, they do not necessarily outperform MBR-Exec with filtering either, suggesting that the information from log-likelihood and “Reviewer” according to Zhang et al. (2022) is not necessarily helping reranking when sufficient direct execution is applied. Note that the MBR-Exec with filtering only slightly underperforms the oracle, suggesting there is little space to propose methods to outperform existing metrics that rely on both filtering and MBR-Exec. Instead, we suggest training a metric that does not rely on execution during inference is a better direction since MBR-Exec is also expensive with our current experimental setting.  [^7] We find difficulties in implementing multi-thread execution to get execution outputs, especially when numerous unit test cases are included (Liu et al., 2023). E.g. CodeLlama-7B-Instruct on HumanEval, generation takes around 3 hours while getting execution outputs takes around 5 hours.
@@ 4.3 Efficient Reranking with Fewer Unit Tests
While MBR-Exec performs well with filtering, it requires access to execution unit test inputs, which is impractical considering the execution cost. Therefore, we ask: how many unit tests should we execute candidates on? With this question in mind, we experiment with 50 generated candidates using MBR-Exec using increasing unit tests. Note that as we use the extended unit tests (Liu et al., 2023), we should also evaluate cases with the extended unit tests. Additionally, when the number of unit tests is lower than those given in the original dataset, we only use unit tests from the original dataset instead of the extended partition. We present results in Figure 5.
@@
Figure 5: Performance of MBR-Exec with fewer unit tests. We only test on HumanEval+ and MBPP+. Note that the original HumanEval benchmark contains an average of 9.6 unit tests per task, and the original MBPP benchmark contains around 3 unit tests per task.
@@
The original HumanEval benchmark contains an average of 9.6 unit tests per task, and the original MBPP benchmark contains around 3 unit tests per task. Therefore, for HumanEval+, we only need 20 unit test cases, while for MBPP-S+, MBR-Exec with 5 unit test cases already gives optimal performance. This suggests that MBR-Exec can already reach its optimal performance with a few unit tests provided. Moreover, we again see the power of bringing in filtering based on trial unit tests, as it substantially improves reranking results across different benchmarks. Specifically, filtering brings a larger impact on HumanEval+ than MBPP-S+, where there’s a larger number of unit tests per task and a lower overlap between trial unit tests and unit tests used for evaluation, suggesting that they potentially have higher testing coverage compared to the trial unit tests provided in MBPP-S+.
@@
Figure 6: Improvement in Pass@k of CodeLlama-7B-Instruct after self-debugging compared to no self-debugging applied.
4.4 Improving the Oracle with Self-Debugging
Model HumanEval(+) MBPP-S(+)
MBR-Exec SD-1 SD-Multi MBR-Exec SD-1 SD-Multi
$\sim$7B Scale
CL-7B-Instruct 52.1 (49.5) 52.0 (49.5) 52.4 (50.0) 64.0 (57.8) 65.8 (59.1) 66.3 (60.4)
   + Filtering 74.5 (70.1) 74.4 (70.1) 76.5 (72.3) 79.0 (69.3) 79.3 (69.2) 81.3 (72.1)
DS-6.7B-Instruct 86.4 (85.5) 87.0 (86.0) 87.7 (87.0) 84.0 (78.6) 84.6 (78.8) 84.9 (79.3)
   + Filtering 91.6 (90.7) 91.5 (90.4) 91.6 (90.4) 89.1 (82.1) 89.0 (81.7) 89.9 (82.8)
$\sim$13B Scale
CL-13B-Instruct 61.3 (59.6) 61.7 (60.1) 62.7 (61.4) 73.9 (67.2) 74.4 (67.0) 74.4 (68.4)
   + Filtering 80.5 (76.1) 80.9 (76.5) 81.6 (77.6) 84.1 (74.7) 83.9 (74.2) 84.9 (76.2)
Table 2: Comparison of self-debugging methods with 50 candidates generated by CodeLlama-{7,13}B-Instruct with temperature 1.6 and DeepSeekCoder-6.7B-Instruct with temperature 1.2, and debugged over one (Chen et al., 2024) or multiple candidates using the same LLM. We also provide the upper bound after debugging. Note that results ending with +++ mean it’s evaluated on the plus with extended test cases. Results are averaged across 4 runs.
Since applying MBR-Exec with filtering based on trial unit tests approaches the theoretical oracle estimated by Pass@k, increasing this oracle seems more sensible than proposing better reranking methods. In this section, we aim to understand the effect of self-debugging on improving the oracle of performance and the performance of reranking generally. We then compare our proposed method of performing SD-Multi to just using one candidate. Note that for SD-Multi, we only use results with one round of debugging, while we use results with 3 rounds of debugging for SD-1.
Results suggest that self-debugging improves oracle performance of reranking, and one single round is enough.
According to Figure 6, we found that self-debugging does help to improve the oracle of reranking, represented by the improvement on Pass@k across different ks. However, later rounds of self-debugging do not show substantial improvement in the oracle compared to the first round.
@@ 4.4 Improving the Oracle with Self-Debugging
According to Table 2, SD-Multi consistently outperforms SD-1 in most experiments, with better-performing models having smaller margins of improvement with self-debugging. The only exception we find is on experiments with candidates generated and self-debugged with DeepSeekCoder-6.7B-Instruct, where we find that the performance of reranking without self-debugging is already over 91, for which we presume that there is little space for improvement.
@@
Figure 7: Comparison of self-debugging methods over different numbers of candidates generated by CodeLlama-7B-Instruct, and debugged over $\{1,\textbf{Multi}\}$ candidates using the same LLM. We also provide the oracle after self-debugging for one round. Note that results ending with $+$ mean it is evaluated on the plus with extended test cases. Results are averaged across at least 4 runs.
Filtering is important in the context of self-debugging.
With 50 candidates generated and self-debugged, more experiments of MBR-Exec outperforming SD-1 with filtering compared to without filtering (see Table 2), suggesting that performance improvement brought by self-debugging can be partly accomplished by filtering, a finding that is missing from Chen et al. (2024) where filtering is also applied but not analyzed. Additionally, when filtering is applied, SD-1 only outperforms MBR-Exec with samples generated and debugged with CodeLlama-13B-Instruct.
The results of Figure 7 show that our proposed method consistently outperforms SD-1 with more rounds of debugging and that with 25 candidates generated, one-time self-debugging is more effective than sampling 50 candidates and debugging on a single candidate obtained by MBR-Exec. The results of MBPP are more moderate than the results of HumanEval, and we acknowledge that filtering closes the gap to the oracle, giving little space for improvement. Last but not least, self-debugging leaves a larger gap between our methods, with state-of-the-art results, and the oracle.
@@ 5 Related Work
Recent years have witnessed the success of pretrained large language models for code generation (Chen et al., 2021; Li et al., 2022, 2023) that have been successful on multiple execution-based benchmarks (Hendrycks et al., 2021; Austin et al., 2021; Chen et al., 2021; Liu et al., 2023; Lai et al., 2022; Wang et al., 2022). More recently, strong-performing open-source pretrained Code LLMs (Rozière et al., 2024; Luo et al., 2024; Guo et al., 2024a; Lozhkov et al., 2024) have been available, and they can perform tasks such as filling-in-the-middle (Rozière et al., 2024), code debugging (Guo et al., 2024b), as some are also trained with instruction tuning. Our works use open-source code LLMs and instruction tuning to study reranking and self-debugging.
Reranking and MBR Decoding.
Reranking and MBR Decoding have been studied extensively in domains such as machine translation (Eikema and Aziz, 2020, 2022; Fernandes et al., 2022; Farinhas et al., 2023). Similarly, Shi et al. (2022) and Li et al. (2022) proposes MBR decoding based on agreement on execution results, with Li et al. (2022) further propose filtering out samples that do not pass the trial unit tests. Shi et al. (2022) propose filtering out samples that cannot get execution outputs, followed by Zhang et al. (2022) who combine this filtering with the log-likelihoods of the generated candidate, and the log-likelihoods of input given the generated candidate. Our work differs from these works as we report the effect of filtering based on trial unit tests on candidate generation, reranking, and self-debugging. Our work also differs in combining reranking with self-debugging.
Other works proposed methods with generated unit test cases with Chen et al. (2023) and To et al. (2024) utilizing generated unit tests for reranking and Huang et al. (2024) incorporating both generated unit test cases and execution-based specifications. Our work differs from this line of research as we adopt the setting with access to trial unit tests and inputs of all unit tests for evaluation to study both filtering based on trial unit tests and MBR-Exec on multiple test cases, while Chen et al. (2023) and To et al. (2024) presume access to neither for prompting or reranking and Huang et al. (2024) only incorporate trial unit tests in some experiments without directly applying filtering based on them. While Chen et al. (2023) experiment with filtering in other algorithmic coding tasks in the appendix, our work differs in showing the effect of filtering on candidate generation, efficient reranking with fewer test cases, and self-debugging. Additionally, our work differs in that we experiment with different candidate generation settings with filtering being applied that leads to findings of new candidate generation “sweet spots”. Our work also differs in combining reranking with self-debugging.
@@
Additionally, previous works have also studied training N-Best rerankers. Early efforts include training classifiers that predict program success in execution (Inala et al., 2022), followed by Ni et al. (2023) proposing to train both model-specific and task-specific classifiers for reranking. Our work differs from Ni et al. (2023) as we evaluate with more general-purpose metrics (Dong et al., 2023) that do not rely on execution during inference time. Moreover, we include filtering based on trial unit tests as part of the feedback, our focus is on comparing execution-based and execution-less metrics. The redundancy of using trained rerankers on target tasks and target code generation models with execution included even during inference time is also supported by our results of filtering with MBR-Exec.
Last but not least, our work differs from all these works as we also study the neural metrics (Zhou et al., 2023; Dong et al., 2023) in comparison to execution-based metrics.
Self-Debugging with LLMs
@@
Chen et al. (2024) study the effect of using the same Code LLM to debug a single generated code candidate by varying feedback information such as execution errors and code explanation. Olausson et al. (2024) also study self-debugging but focus on different sampling methods and their impacts on self-debugging. Zhong et al. (2024) additionally integrate information regarding intermediate states to provide additional information for debugging. Our work differs from theirs as we study the joint effect of both self-debugging and reranking. We also perform self-debugging on multiple candidates.

### v4

@@ 4.1 Candidate Generation
(c) LiveCodeBench
Figure 2: Performance of reranking and oracle performance over different numbers of generated candidates using CodeLlama-7B-Instruct with temperature 1.6 for HumanEval+ and MBPP+, and 1.2 for LiveCodeBench. Results are averaged across at least 2 runs for LiveCodeBench and 4 runs for the rest.
What is the impact of the number of candidates?
We generate candidates using CodeLlama-7B-Instruct by setting temperature 1.6 and p 0.95.
Figure 2 shows the oracle performance and the reranking performance. First of all, when filtering is not included, MBR-Exec only shows moderate improvement with more generated candidates across different numbers of unit tests for evaluation and two benchmarks, aligning with the analysis from Shi et al. (2022). However, with filtering, MBR-Exec shows considerable improvements over the case without filtering, with performance scaling much better with more candidates generated, which aligns with our hypothesis. Secondly, the performance of MBR-Exec with filtering consistently approaches the oracle performance even as we increase the number of candidates, providing further evidence supporting our hypothesis.
What is the impact of sampling temperature?
We compare generations with temperatures using CodeLlama-7B-Instruct with nucleus $p$ 0.95 50 generated candidates. We show results in Figure 3.
@@
(c) LiveCodeBench
Figure 3: Performance of reranking and oracle over sampling temperatures using CodeLlama-7B-Instruct with 50 generated candidates over 4 runs.
We find that sampling with a previously unseen high gives a higher oracle of reranking performance, peaking between 1.6 and 1.8 for HumanEval+ and MBPP+, and between 1.2 and 1.4 for LiveCodeBench. This suggests that sampling with a lower temperature is suggested when applied to tasks that are more difficult, e.g. competitive programming when compared to basic programming. However, sampling in practice generally can be done with a higher temperature than previously suggested (Chen et al., 2021; Li et al., 2022; Shi et al., 2022; Liu et al., 2023; Jain et al., 2024). Additionally, when filtering is applied, reranking performance approaches the oracle, which again supports our aforementioned hypothesis.
For later experiments, we choose the number of generated candidates to be 50. We adopt sampling temperature 1.6 for HumanEval and MBPP-S and 1.2 for LiveCodeBench, with nucleus $p$ 0.95. For DeepSeek models, we use temperature 1.2333With temperature ¿1.2 on DeepSeek models, sampling is more likely to generate token indices that are not defined in the vocabulary, see https://github.com/vllm-project/vllm/pull/3685..
HE+ MBPP-S+ LCB
Random 30.2 41.9 13.3
Greedy 39.0 44.8 13.4
Oracle 76.2 72.0 44.1
@@
LL 38.7 42.6 16.3
CR 40.5 43.1 15.9
CS 30.2 - 14.1
Filter 59.7 60.8 29.6
Filter + LL 60.4 59.7 32.5
Filter + CR 61.3 60.2 32.2
Filter + CS 61.0 - 30.2
@@
LL + MBR-Exec 49.4 58.1 24.2
CR + MBR-Exec 48.9 58.7 24.2
Filter + MBR-CBS 60.1 61.3 -
Filter + MBR-CS 59.9 - -
Filter + MBR-Exec 70.1 69.3 38.7
Filter + LL + MBR-Exec 70.6 69.2 33.4
Filter + CR + MBR-Exec 70.0 69.4 33.2
Table 1: Comparison of reranking methods on HumanEval (HE)+, MBPP-S+, and LiveCodeBench(LCB). For N-Best Reranking, we compare Likelihood (LL), Coder-Reviewer (CR), and CodeScore (CS). For MBR, we compare CodeScore, CodeBertScore (CBS), and MBR-Exec. We highlight best and second best reranking results of the class of reranking methods.
@@ 4.2 Comparing Reranking Methods
We conduct experiments with $n$-best reranking and MBR to compare different raking approaches with and without filtering based on trial unit tests. We show results in Table 4.
Filtering is highlighted.
According to the group “N-Best Reranking” (see Table 4), filtering based on trial unit tests gives the largest gain from the random baseline compared to the likelihood feature, CoderReviewer, and CodeScore.
Not only is filtering powerful in itself, but it also shows further improvements when combined with $n$-best reranking methods. Therefore, we again emphasize the importance of filtering based on unit tests for reranking.
@@
Zhang et al. (2023) shows that Coder-Reviewer outperforms MBR-Exec with a weaker type of filtering, i.e., retaining a code candidate if it can produce outputs given the inputs provided in trial unit tests. In our experimental settings that include a more sensible definition of filtering and assume access to all unit test inputs for evaluation, likelihood feature and CoderReviewer underperform MBR-Exec with filtering. Additionally, when combined with filtering and MBR-Exec, they do not substantially outperform MBR-Exec with filtering either, suggesting that the likelihood feature and CoderReviewer (Zhang et al., 2023) do not substantially help reranking when sufficient direct execution is applied, giving close-to-oracle performances.
Additionally, when combined with filtering, MBR based on neural metrics does not significantly outperform filtering itself on basic programming tasks (HumanEval and MBPP). Therefore, we do not recommend compared to using execution-based MBR. Acknowleding the huge computational cost of applying neural metrics for MBR, we do not apply it to competitive programming with long generated sequences.
@@ 4.3 Efficient Reranking with Fewer Unit Tests
While MBR-Exec performs well with filtering, it requires access to execution unit test inputs, which is impractical considering the execution cost. Therefore, we ask: how many high-quality unit tests should we execute candidates on? We experiment with 50 generated candidates using MBR-Exec using increasing unit tests. Note that as we use the extended unit tests (Liu et al., 2023), we should also evaluate cases with the extended unit tests. We do not test this on LiveCodeBench due to the limit of unit tests. Additionally, when the number of unit tests is lower than those given in the original dataset, we only use unit tests from the original dataset instead of the extended partition.
@@
Figure 4: Performance of MBR-Exec with fewer unit tests.
@@
According to Figure 4, for HumanEval+, we only need 20 unit tests, while for MBPP-S+, MBR-Exec with 5 unit tests already gives optimal performance. This suggests that MBR-Exec can already reach its optimal performance with a few unit tests provided.
4.4 Improving Oracle with Self-Debug
Model HumanEval+ MBPP-S+ LiveCodeBench
MBR SD-1 SD-Multi MBR SD-1 SD-Multi MBR SD-1 SD-Multi
$\sim$7B Scale
CL-7B-Instruct 49.5 49.5 50.0 57.8 59.1 60.4 23.7 23.3 24.0
   + Filtering 70.1 70.1 72.3 69.3 69.2 72.1 38.7 38.0 39.9
DS-6.7B-Instruct 85.5 86.0 87.0 78.6 78.8 79.3 30.6 30.6 31.3
   + Filtering 90.7 90.4 90.4 82.1 81.7 82.8 41.9 41.2 42.0
$\sim$13B Scale
CL-13B-Instruct 59.6 60.1 61.4 67.2 67.0 68.4 27.0 27.6 27.6
   + Filtering 76.1 76.5 77.6 74.7 74.2 76.2 40.6 40.8 46.8
$\sim$16B Scale
DS-V2-Lite-Instruct 83.5 83.4 85.1 80.1 81.0 81.4 45.8 45.4 46.6
   + Filtering 89.6 88.4 89.6 81.8 82.2 82.8 58.7 58.3 58.5
Table 2: Comparison of self-debugging methods with 50 candidates generated by CodeLlama-{7,13}B-Instruct and DeepSeekCoder-{6.7B,V2-Lite}-Instruct, and debugged over {1, Multi} candidates. We also provide the upper bound after debugging. Results are averaged across 2 runs for LiveCodeBench and 4 runs for the rest.
Applying MBR-Exec with filtering gives close-to-oracle performances, suggesting that improving the oracle performance with Self-Debug (Chen et al., 2024) is a more sensible choice than providing better reranking methods. In this section, we first analyze the improvement in oracle performance. We then compare our proposed SD-Multi and SD-1 proposed by Chen et al. (2024). For SD-Multi, we only use results with one round of debugging, while we use results with 3 rounds of debugging for SD-1.
Self-Debug improves oracle performance of reranking, and one single round is enough.
According to Figure 5, Self-Debug helps improve the oracle of reranking, represented by the improvement on Pass@k across different ks. However, later rounds of Self-Debug do not show substantial improvement in the oracle compared to the first round. Moreover, Self-Debug with one single round LiveCodeBench shows larger Pass@k improvement with larger ks. A sensible guess for this trend is related to the higher level of difficulty of problems in LiveCodeBench, with Self-Debug regenerating few sensible programs on problems that originally had no sensible candidates. The disappearance of this improvement suggests that since most code LLMs have not been trained extensively on debugging objectives, Self-Debug regenerates candidates close to greedy decoding.
@@ 4.4 Improving Oracle with Self-Debug
(c) LiveCodeBench
Figure 5: Improvement in Pass@k of CodeLlama-7B-Instruct after Self-Debug compared to no Self-Debug applied.
@@
According to Table 2, SD-Multi consistently outperforms SD-1 in most experiments, with better-performing models having smaller margins of improvement with Self-Debug. The only exceptions we find are on experiments with candidates generated and self-debugged with DeepSeekCoder-{6.7B, V2-Lite}-Instruct, where either the task is considered more difficult compared to basic programming, or that the baseline performance with MBR-Exec is already above 91.
@@
(c) LiveCodeBench
Figure 6: Comparison of Self-Debug methods over different numbers of candidates generated by CodeLlama-7B-Instruct, and debugged over $\{1,\textbf{Multi}\}$ candidates. Results are averaged across at least 4 runs (and 2 runs for LiveCodeBench).
Filtering matters in Self-Debug.
With 50 candidates generated and self-debugged, more experiments of MBR-Exec outperforming SD-1 with filtering compared to without filtering (see Table 2), suggesting that improvement brought by Self-Debug can be partly accomplished by filtering, a finding that is missing from Chen et al. (2024) where filtering is also applied but not analyzed. Additionally, when filtering is applied, SD-1 only outperforms MBR-Exec with samples generated and debugged with CodeLlama-13B-Instruct.
Figure 6 shows that SD-Multi (1 round) consistently outperforms SD-1 (3 rounds) with different numbers of candidates generated. Last but not least, Self-Debug leaves a larger gap between SD-Multi and the oracle.
@@ 5 Related Work
Recent years have witnessed the success of pretrained large language models for code generation (Chen et al., 2021; Li et al., 2022, 2023; Sun et al., 2024; Zhuo, 2024) that have been successful on multiple execution-based benchmarks (Hendrycks et al., 2021; Austin et al., 2021; Chen et al., 2021; Liu et al., 2023; Lai et al., 2023; Wang et al., 2023b). More recently, strong-performing open-source pretrained Code LLMs (Rozière et al., 2023; Luo et al., 2024; Guo et al., 2024a; Lozhkov et al., 2024; DeepSeek-AI et al., 2024) have been available, and they can perform tasks such as filling-in-the-middle (Rozière et al., 2023), code debugging (Guo et al., 2024b). Our work uses representative open-source code LLMs to study generation, reranking, and Self-Debug.
Reranking in Code Generation
Shi et al. (2022) and Li et al. (2022) proposed MBR decoding based on agreement on execution results. Li et al. (2022) further proposed filtering based on trial unit tests, and Shi et al. (2022) proposed filtering out samples that cannot get execution outputs, followed by Zhang et al. (2023) combining log-likelihoods of the code candidate and the instruction. Farinhas et al. (2024) further studied the reranking laws based on communication theory. Our work systematically studies proposed methods.
Other works adopt a setting without unit test inputs for evaluation available, and proposed methods with generated unit test cases utilizing generated unit tests (Li et al., 2022; Chen et al., 2023; To et al., 2024) for reranking and Huang et al. (2024) incorporating both generated unit test cases and functional specifications Our work differs from these efforts as we adopt a setting with unit test inputs for evaluation available.
@@
Additionally, previous works have also studied training N-Best rerankers. Early efforts include training classifiers that predict program success in execution (Inala et al., 2022), followed by Ni et al. (2023) proposing to train both model-specific and task-specific classifiers for reranking. Our work differs from Ni et al. (2023) as we evaluate with more general-purpose metrics (Dong et al., 2023) that do not rely on execution during inference time.
Generally, our work differs as we systematically study the reranking with representative methods, along with sampling strategies and Self-Debug.
Self-Debug with LLMs
@@
Chen et al. (2024) studied the effect of using the same Code LLM to debug a single generated code candidate by varying feedback information such as execution errors and code explanation. Olausson et al. (2024) also studied Self-Debug but focus on different sampling methods and their impacts on Self-Debug. Zhong et al. (2024) additionally integrated information regarding intermediate states to provide additional information for debugging. Our work differs from theirs as we study the joint effect of both Self-Debug and reranking. We also perform Self-Debug on multiple candidates.